The new tires in George’s vehicle are 15.5 inches in diameter. They are guaranteed for 75,000 miles, meaning that the thread will wear less than 0.50 inches in 75,000 miles. When he got his new tires, George had determined that he was getting an average of 34.55 miles per gallon of fuel. He got suspicious of the tires when, after 25,000 of driving, his mileage improved to 35.00 miles per gallon; and he was sure of his problem when after another 25,000 miles, his mileage was 35.47 miles per gallon.
Please, use the information given to help George make a case to return his defective tires.

Answer by ankor@dixie-net.com(12702)

The new tires in George’s vehicle are 15.5 inches in diameter. They are guaranteed for 75,000 miles, meaning that the thread will wear less than 0.50 inches in 75,000 miles. When he got his new tires, George had determined that he was getting an average of 34.55 miles per gallon of fuel. He got suspicious of the tires when, after 25,000 of driving, his mileage improved to 35.00 miles per gallon; and he was sure of his problem when after another 25,000 miles, his mileage was 35.47 miles per gallon.
Use the information given to help George make a case to return his defective tires.
:
I guess they mean that the recorded mileage is based on the number of revolutions of a new tire's circumference.
:
Assuming he is using the odometer on the car, the number of revolutions per gallon of gas will increase, (falsely) indicating better gas mileage.
:
Using the difference in circumference:
15.5/15.0 * 34.55 = 35.7 mpg expected at 75,000 miles
:
After 25000 mi, 1/3 of the life of the tire. 2/3 of the tire life remaining
2/3 * 1/2 = .33333; Circumference should be 15.3333 in
15.5/15.33333 * 34.55 = 34.92 mpg expected
:
Another 25000 mi, 2/3 of the life of the tire. 1/3 of the tire life remaining
1/3 * 1/2 = .1666667, Circumference should be 15.166667 in
15.5/15.166667 * 34.55 = 35.31 mpg expected
:
I don't think you have enough difference to make a case. Many other factors
such as driving conditions, driving habits, season of the year, brand of gas etc :
Expected | Actual
Start 34.55 | 34.55
25k: 34.92 | 35.00
50k: 35.31 | 35.47
: